OpenStudy (anonymous):

Help :( I know the answer but not how to get there. find indicated derivative d/du ((u)/(u-1))-((u)/(u+1))

OpenStudy (anonymous):

Do you know the quotient rule?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Just take the derivative of each fraction using the quotient rule and combine them

OpenStudy (anonymous):

i get \[\frac{( u-1)-u }{ u-1^{2} } - \frac{( u+1)-u }{ u+1^2 }\]

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

Ok so let's just say that the top is x and the denominator is y... x/y the quotient rule would be: \[ \frac{ (x' * y)-(y' * x)}{ y ^{2}}\]

OpenStudy (anonymous):

yeah i tried the quotient rule for both and got that, not combined yet, i think i did it wrong though

OpenStudy (anonymous):

The derivative of the first fraction would be: \[ \frac{ (1*(x-1)) - (1*x)}{ (x-1)^{2}}\]

OpenStudy (anonymous):

It looks like you just forgot to square the entire denominator

OpenStudy (anonymous):

okay, i got that. now the second one right?

OpenStudy (anonymous):

The second on looks the same way. Just make sure you square the entire denominator not just part of it!

OpenStudy (anonymous):

i just forgot to put the parenthesis on the typing part lol, but i dont know how my math textbook got -[((u)/((u-1)^2) - ((u)/(u+1)^2)]

OpenStudy (anonymous):

as the answer

OpenStudy (anonymous):

hmm... that's not right! The numerators for both would be 1! I double checked it too! http://www.wolframalpha.com/input/?i=d%2Fdu+%28%28u%29%2F%28u-1%29%29-%28%28u%29%2F%28u%2B1%29%29

OpenStudy (anonymous):

see, i hate when textbooks have typos :/ they make me think i did something wrong lmao. thanks

OpenStudy (anonymous):

no problem!